MSU linebacker Denicos Allen sacks UM quarterback Denard Robinson. Just after they collide, they are momentarily stuck together, and both are moving at a speed of 2.80 m/s. Denicos Allen's mass is 101.1 kg and Denard Robinson's mass is 92.0 kg. If the quarterback was at rest just before he was sacked, how fast was the linebacker moving just before the collision?
(in m/s)
Answer #1

m1u1 + 0 = (m1+m2)V

m1u1 = 193.1*2.8

u1 = 540.68 /101.1 = 5.34 m/s